OpenStudy (anonymous):

During a race, a runner runs at a speed of 6 m/s. 2 seconds later, she is running at a speed of 10 m/s. What is the runner’s acceleration. also can you show ur work so i understand

OpenStudy (deadly_roses):

a = (v2 - v1)/t = (10 - 6)/2 = 2 m/sec/sec (average acceleration)

OpenStudy (anonymous):

thanks bae

OpenStudy (deadly_roses):

no problem :)